OpenStudy (anonymous):

Solve the equation using the Zero-Product Property. (x-2)(x+7)=0 A. 2, -7 B. 2, 7 C. –1,1 D. –2, -7

OpenStudy (phi):

what is the zero-product property?

OpenStudy (anonymous):

is the answer A?

OpenStudy (phi):

A says x can be 2 or -7 if x were 2 in (x-2)(x+7)=0 you replace the x with 2: (2-2)(2+7)=0 2-2 is 0 and 0 times anything is 0. so 2 works if x were -7: (-7 -2 )(-7+7) =0 -7+7 is 0, and 0 times -9 is 0. -7 works